OpenStudy (anonymous):

michael takes 90 min to type 5850 word if he increases his typing speed by 5 word /mins ,how many minutes will he take to type 10,500 words

OpenStudy (ja1):

you want the answer or to work it out?

OpenStudy (anonymous):

work it out please because i want to see where i went wrong

OpenStudy (ja1):

alright brb

OpenStudy (ja1):

Here is the starting equation 5850/90=65

OpenStudy (anonymous):

ok i got that right

OpenStudy (ja1):

now 65+5=70

OpenStudy (anonymous):

ok that where i messed up i subtracted instead of add but thanks keep going

OpenStudy (ja1):

oh ok now 70x=10500

OpenStudy (anonymous):

ok then divide both sides by 70 right an i got 150

OpenStudy (ja1):

yupp :D glad i helped

OpenStudy (anonymous):

can u help me with another problem i realy dont understand it

OpenStudy (ja1):

sure

OpenStudy (anonymous):

tara works in a clothing store where she earns a base salary of 100 per day plus 12% of her daily sales .she sold 800 in clothing on sunday .how much did she earn over the two days

OpenStudy (ja1):

ok so im guessing thatyour going to include the 12% only for sunday then 100 for the next day right?

OpenStudy (ja1):

the problem itself is easy but the wording is hard, why does it say over the two days when it only said sunday?

OpenStudy (anonymous):

yes thats the only thing i did so far

OpenStudy (ja1):

still dont know why it says how much did she earn over the two days

OpenStudy (ja1):

but either way lets go

OpenStudy (ja1):

so first what is 12% of 800

OpenStudy (anonymous):

96

OpenStudy (ja1):

yes, now if im right then its 100 + 100 + 96

OpenStudy (ja1):

which is?

OpenStudy (anonymous):

296 the answer choices are 276 376 476 and 576

OpenStudy (ja1):

?

OpenStudy (ja1):

ok just a sec

OpenStudy (anonymous):

ok

OpenStudy (ja1):

what class is this algebra 1?

OpenStudy (anonymous):

yes algebra 1 honors

OpenStudy (ja1):

7th grade right

OpenStudy (anonymous):

no 9th

OpenStudy (ja1):

oh thought you were doing 7th grade but accelerated course, anyways just a sec

OpenStudy (anonymous):

ok

OpenStudy (ja1):

you sure the problem is written correctly, did she sell any clothes on saturday?

OpenStudy (anonymous):

im so sorry 800 on saturday n 1500 sunday

OpenStudy (ja1):

oh lol that makes more sense

OpenStudy (ja1):

so she made 196 on saturday since she sold 800 and on sunday she sold 100 + (12% of 1500)

OpenStudy (anonymous):

280

OpenStudy (ja1):

yes so 280 + 196 = answer :D easy

OpenStudy (anonymous):

476

OpenStudy (ja1):

yupp thats it, see not so hard its just the wording that makes it hard

OpenStudy (anonymous):

yess no i see

OpenStudy (ja1):

yay, well im glad i helped, you mind giving me a medal :)

OpenStudy (ja1):

thanks :)

OpenStudy (anonymous):

your welcome